# PhishLumos: An Adaptive Multi-Agent System for Proactive Phishing Campaign Mitigation

### v1

@@ I Introduction
Phishing attacks, which use social engineering to steal confidential data, represent a persistent and adaptable security threat. The impact is substantial: total losses from cybercrime in the U.S. exceeded \$12.5 billion in 2023 alone. Within this landscape, phishing stands out as the most frequently reported type of crime, highlighting its central role as a societal threat [1]. These attacks deepen the digital divide by disproportionately harming vulnerable populations, including older adults and those with less technical expertise [2, 3, 4]. Critically, this erosion of trust in essential digital services—from banking to healthcare—threatens not just economic stability but also digital equity in our increasingly online society. Defensive strategies have long relied on tools like blocklists, which are typically populated based on data from automated security scanners. However, today’s attackers employ sophisticated evasive tactics, such as cloaking (delivering different content to scanners versus users), which makes these established methods less effective. This situation demands more advanced, proactive, and resilient solutions to counter these modern threats.
@@
Figure 1: The core concept of PhishLumos. While conventional security scanners fail when faced with cloaking, PhishLumos treats such evasion as a key signal to pivot its investigation toward peripheral infrastructure footprints, allowing it to uncover the entire malicious campaign.
@@ II Related Work
Detecting advanced phishing campaigns remains a significant challenge, especially when attackers use evasive tactics that limit available data. Previous research has explored several methods, including content-based analysis, countermeasures for cloaking, and the use of LLMs. However, these approaches generally assume that enough webpage content is available for analysis. They also typically focus on identifying individual malicious URLs. They do not specialize in uncovering entire campaigns or creating detection rules from limited initial clues. PhishLumos addresses this specific gap. This section reviews prior work and explains how PhishLumos provides adaptive, campaign-level detection, which is effective even when starting with minimal information.
Content-Based Phishing Detection. Content-based detection is a common defense against phishing, but it becomes less effective when content is hidden or minimal. PhishLumos is designed specifically for this scenario. Methods like VisualPhishNet [10] (visual similarity), Phishpedia [11] (logo detection), and PhishIntention [12] (intent identification) are powerful. However, their success depends on having rich and analyzable webpage data. Consequently, they struggle in content-inaccessible environments featuring cloaking or minimal content. PhishLumos is specifically designed to work effectively in these scenarios. It uses peripheral infrastructure data to uncover entire campaigns, a capability that goes beyond direct content inspection.
Evasion Techniques and Countermeasures. Prior research has thoroughly documented evasion techniques and the efforts to counter them. Early work by Invernizzi et al. explored web cloaking [13], and CrawlPhish detailed client-side evasions [6]. Studies like PhishFarm [7] and PhishTime [14] confirmed that cloaking is effective. Countermeasures include PhishDecloaker [5] for bypassing CAPTCHAs, PhishPrint [8] for analyzing how scanners are exploited, and even offensive systems like Spartacus [15]. While these approaches focus on bypassing or understanding evasion, PhishLumos uniquely uses the inability to access content as the primary signal to initiate its infrastructure investigation. This signal becomes the starting point for generating campaign-level detection rules.
Multimodal Approaches and LLM Utilization. The use of LLMs and multimodal data has led to powerful new methods for phishing detection. However, most of these systems still rely on having accessible webpage information. For instance, KnowPhish [16] employs LLMs to improve detection based on reference pages. PhishLLM [9] and ChatPhishDetector [17] use the reasoning capabilities of LLMs to analyze crawled web data, such as HTML, text, and screenshots, to find malicious intent. Other systems, like ScamNet [18] and PhishAgent [19], use specialized LLMs for fraud detection. The primary difference from these approaches, which excel with rich content, is that PhishLumos is optimized for content-inaccessible scenarios. Instead of simply analyzing given content, its LLM agents are designed to initiate and direct an investigation using minimal data. PhishLumos uses LLMs to dynamically search external data sources, connect separate clues to build a campaign profile, and create practical detection rules. This focus on adaptive, LLM-driven investigation from scarce information sets it apart from previous work.
@@ III Proposed System
Agent Primary Goal Key Inputs & Outputs
Supervisor Orchestrate overall analysis flow. Input: URL. Output: Final Report.
Specialized Agents
URL Analysis Get historical scan data for a URL. Input: URL. Output: Scan history, status, IP addresses, brands.
Domain Analysis Analyze historical DNS/IP data. Input: Domain Name. Output: IP address history, stability patterns.
IP Analysis Analyze domains/activity on an IP. Input: IP Address. Output: Associated domains, past malicious activity.
Certificate Analysis Analyze TLS certificate history. Input: Domain Name. Output: Issuer patterns, shared certificates.
Synthesis Agents
Campaign Analysis Synthesize all findings to profile a campaign. Input: All collected data. Output: Campaign type, key indicators.
Rule Generation Generate and validate detection rules. Input: Key indicators. Output: Validated search queries.
III-A System Overview
We introduce PhishLumos, an adaptive multi-agent system for detecting advanced phishing attacks, particularly those employing evasive techniques that limit content availability. Figure 2 illustrates the system architecture. Unlike traditional methods that struggle with insufficient data, PhishLumos is designed for such content-inaccessible scenarios. It performs what we term Pivotal Analysis: an adaptive process where LLM-powered agents shift the investigative focus from an evasive URL to its underlying infrastructure footprints. This involves targeted analysis of various external datasets, including historical DNS records, IP address information, and TLS certificate data.
Figure 2: A Supervisor Agent dynamically tasks Specialized Agents, whose findings update a shared Knowledge Base. This evolving context guides the Pivotal Analysis loop. Finally, Synthesis Agents analyze the full Knowledge Base to create a campaign profile and validated detection rules.
@@ III-B Threat Model
Social Dimension. These attacks erode digital equity by targeting vulnerable populations, such as older adults. Beyond financial loss, they shatter the trust required to use essential online services, risking social isolation and loss of autonomy [2, 20].
Technical Dimension. We assume attackers use widely available phishing kits. A key feature of these kits is sophisticated cloaking logic, which delivers malicious content to end-users while presenting benign content to automated scanners. We consider email and SMS as primary distribution channels. The analytical focus of PhishLumos is on characterizing and mitigating these coordinated campaigns that deploy many similar, cloaked sites concurrently, rather than isolated, one-off attacks.
@@ III-D Adaptive Pivotal Analysis
At the core of the loop, the LLM assesses the entire Knowledge Base to identify the most promising investigative leads based on a set of triggers. These include: (1) Immediate Triggers, such as significant cloaking behavior (e.g., a newly registered domain redirecting to an unrelated major site like google.com); and (2) Infrastructure Triggers, where accumulated evidence reveals suspicious hosting patterns, such as multiple unrelated domains on a non-CDN IP address. This state-aware, trigger-based orchestration allows PhishLumos to efficiently allocate analytical resources, adapting its focus to the specific tactics employed by each unique campaign.
@@ III-E Data Acquisition and Context Management
PhishLumos employs a systematic process for data acquisition and context management to support its adaptive investigations. External intelligence is sourced programmatically from threat intelligence APIs, such as urlscan.io [21] and VirusTotal [22], which provide access to essential datasets such as historical website scans, DNS records, and TLS certificate attributes. Specialized Agents dynamically retrieve this information using dedicated API interaction tools.
@@ III-F Campaign Analysis and Rule Generation
TABLE II: Primary evidence indicators for campaign classification by the Campaign Analysis Agent.
@@
CONFIRMED Consistent brand detection or content hashes,
linked by shared IP addresses.
CLOAKED Discrepancy between input and final URL,
especially to a known legitimate site.
REUSE Shared IPs/Certificates with historically
malicious domains (content may differ).
UNAVAILABLE Inaccessible content (e.g., 404), but
infrastructure linked to a known campaign.
UNCLEAR Insufficient correlated evidence to make
a confident classification.
@@
III-G Illustrative Example
To demonstrate the adaptive workflow of PhishLumos, we present a sanitized example based on a real-world phishing campaign. The system receives a single URL, https://portal-service.example.test, which initially shows no direct signs of being malicious.
1) Initial URL Analysis: The URL Analysis Agent queries a web analysis service for portal-service.example.test. The results show a 200 OK status, but no brand impersonation is detected. The page resolves to IP address 203.0.113.55 and uses a TLS certificate issued by a generic provider, Generic-Cert-Issuer. Lacking direct evidence, the Supervisor pivots to infrastructure analysis.
2) IP Analysis Pivot: The Supervisor tasks the IP Analysis Agent with investigating 203.0.113.55. This is a critical pivot. The agent finds this IP has hosted other domains. Crucially, one of these historical domains, secure-auth.example.org, was associated with a detected brand impersonation of Service-A just weeks prior. This finding is the first major breakthrough, linking the non-descript input URL to confirmed malicious activity via shared infrastructure.
3) Certificate Analysis Pivot: Concurrently, the Certificate Analysis Agent examines the certificate associated with the hosting infrastructure. It confirms the generic issuer and notes the certificate’s short validity, a common phishing tactic.
4) Campaign Analysis Synthesis: The Campaign Analysis Agent receives all the collected evidence. The most critical piece is the link established by the IP Analysis Agent: the input URL’s IP (203.0.113.55) was recently used to host a confirmed phishing site (secure-auth.example.org) targeting Service-A. Although the current URL shows no brand, the shared infrastructure provides a strong link. The agent classifies this as an REUSE campaign with HIGH confidence.
5) Rule Generation and Evaluation: The Rule Generation Agent receives the key indicators, primarily the shared IP address (203.0.113.55) and the historically associated detected brand (Service-A). It then formulates a logical detection rule. For this campaign, the generated rule flags a URL if its final landing IP address is 203.0.113.55, and if at least one of the following is also true: the system detects brand impersonation of Service-A, or the scan is for a domain on that IP within a specific date range. This rule is then used to construct a validated query that successfully identifies both current and past sites, effectively mapping the campaign’s footprint.
This example illustrates how PhishLumos uses Pivotal Analysis to uncover connections that are invisible from the initial URL alone. By treating the IP address as a pivot point, the system identified a malicious campaign from seemingly benign evidence.
@@ IV-A Experimental Setup
Datasets. To evaluate PhishLumos, we constructed two primary datasets: a ground truth collection of phishing URLs and a set of benign URLs for false positive analysis. Our ground truth phishing dataset was built using public data from the Japan Computer Emergency Response Team Coordination Center (JPCERT/CC) [23], spanning from January 2023 to March 2025. JPCERT/CC is Japan’s national Cyber Security Incident Response Team (CSIRT), and its data is highly reliable as human analysts manually verify each reported phishing URL. We identified and validated 103 distinct campaigns, comprising 6,020 unique URLs. For false positive analysis, we used a benign dataset of 1,000 safe URLs from the ChatPhishDetector study [17], including legitimate corporate and top-ranking websites.
The Challenge of Content-Inaccessible Scenarios. A critical challenge in modern phishing detection is the prevalence of evasive tactics. Our analysis of the ground truth URLs revealed that automated web analysis APIs often fail to retrieve direct phishing content. Cloaking or access errors affected 77.0% of URLs (82.5% of campaigns), preventing direct content retrieval. This prevalence of content-inaccessible scenarios underscores the limitations of content-based detectors and highlights the need for infrastructure-focused analysis like PhishLumos. A breakdown reveals diverse evasive tactics: 38.8% used cloaking or deceptive redirection, 37.6% were inaccessible or actively blocked scanners (e.g., with 403/404 errors), and the remaining 23.5% presented ambiguous content with no clear malicious indicators. The problem is escalating, with the rate of these content-inaccessible scenarios increasing from 48.7% in 2023 to 95.4% in 2024, and reaching 100% for the campaigns observed in early 2025.
@@
Implementation and Reproducibility. PhishLumos is implemented in Python 3.11.12, using the LangGraph library (v0.4.5) to manage the multi-agent workflow. All agents use Azure OpenAI’s gpt-4.1-2025-04-14 model. To ensure deterministic outputs and maximize reproducibility, key hyperparameters were set to fixed values. We set a temperature of 0 and a seed of 42 for this purpose. Given the deterministic nature of this setup, each experiment was conducted once per input URL. For external data gathering, the agents interfaced with the public API of urlscan.io [21]. All experiments were run on a MacBook Air with an Apple M2 CPU and 24GB of RAM, running macOS Sequoia 15.5.
TABLE III: Overall performance of PhishLumos on proactive campaign mitigation across 103 campaigns.
Metric Mean Median Total
Effectiveness
Campaign Coverage (%) 93.0 100.0 -
New URLs Found 751.4 297.0 77,391
Social Impact
Detection Lead Time (hrs) 529.1 192.8 -
Efficiency
Execution Time (sec) / URL 63.9 60.4 -
Cost (\$) / URL 0.25 0.23 \$26.26
TABLE IV: Performance comparison in different scenarios. PhishLumos excels where content-based methods struggle.
Scenario Method Prec. Recall F1 FPR
@@
TABLE V: Representative case studies demonstrating PhishLumos’s advanced reasoning. Note: Brand names and technical indicators are anonymized for security while preserving the core characteristics of the original campaigns. Initial refers to the properties of the input URL (such as the URL path or its domain), while Final refers to the page attributes after any redirects.
Case Study Challenge Generated Rule Insight: Why this rule is intelligent and practical
1. Financial A Zero initial Initial Domain: *.co.jp.*.test AND Pivoted from a dead-end URL to its domain, spotting
scan data Date:[2023-09-02 TO 2024-02-27] a suspicious domain pattern. This uncovered a regional
(dead-end URL) campaign, finding 405 new URLs.
2. Financial B Deceptive cloaking Initial URL: */[unique_string]* AND Identified a non-obvious URL path string as a unique
(redirect to benign Date:[2024-11-26 TO 2024-12-24] campaign fingerprint, ignoring the misleading redirect.
site) Discovered 540 new URLs.
3. Transportation C Advanced cloaking Initial Domain: *.suspicious.test AND Defeated cloaking by flagging the mismatch between
(redirect to official Final Domain: official.example AND the suspicious initial and legitimate final domains.
site) Date:[2023-04-24 TO 2023-05-15] Found 281 new cloaked URLs.
4. E-commerce D Scanner blocking Final IP: 203.0.113.43 AND Turned a ‘403 Forbidden’ error into a detection rule,
(403 Forbidden Final Status: 403 AND treating the evasion tactic itself as a cloaking signal.
error) Date:[2024-03-30 TO 2024-07-26] Identified 65 new URLs.
5. E-commerce E IP hopping to Final ASN: AS64500 AND Created a resilient rule by targeting the stable ASN
evade blocklists. Final Domain: *.test AND instead of transient IPs, moving up the network stack.
Date:[2023-06-14 TO 2024-01-01] Preemptively found 4,118 URLs.
TABLE VI: Ablation study results. Performance is measured by campaign coverage (%).

### v2

@@ I Introduction
Phishing attacks, which use social engineering to steal confidential data, represent a persistent and adaptable security threat. The impact is substantial: total losses from cybercrime in the U.S. exceeded \$12.5 billion in 2023 alone. Within this landscape, phishing stands out as the most frequently reported type of crime, highlighting its central role as a societal threat [1]. These attacks deepen the digital divide by disproportionately harming vulnerable populations, including older adults and those with less technical expertise [2]. Critically, this erosion of trust in essential digital services—from banking to healthcare—threatens not just economic stability but also digital equity in our increasingly online society. Defensive strategies have long relied on tools like blocklists, which are typically populated based on data from automated security scanners. However, today’s attackers employ sophisticated evasive tactics, such as cloaking (delivering different content to scanners versus users), which makes these established methods less effective. This situation demands more advanced, proactive, and resilient solutions to counter these modern threats.
@@
Figure 1: The core concept of PhishLumos.
@@ II Related Work
Detecting coordinated phishing campaigns is difficult when evasion hides page content from automated scanners. Prior work spans (1) content-based detectors, (2) measurements and countermeasures for cloaking, and (3) LLM and multimodal methods. Most assume accessible content and target single URLs, leaving little support for campaign discovery from sparse signals. PhishLumos addresses this gap by starting from minimal evidence and producing campaign-level rules.
Content-Based Phishing Detection. Visual similarity, logo recognition, and intent inference drive recent detectors such as VisualPhishNet [8], Phishpedia [9], and PhishIntention [10]. These systems are effective when pages expose text, layout, or images, but performance degrades when content is hidden or unavailable. Unlike these methods, PhishLumos operates in content-inaccessible cases (scanners cannot retrieve meaningful content) by pivoting to infrastructure signals and grouping pages into campaigns.
Evasion Techniques and Countermeasures. Work on cloaking and client-side evasions includes CrawlPhish [4] and Invernizzi et al. [11]; controlled measurements in PhishFarm [5] and PhishTime [12] showed their effectiveness. Tools such as PhishDecloaker [3], PhishPrint [6], and PhishParrot [13] aim to bypass or analyze these tactics. PhishLumos differs by treating inaccessibility as a trigger to investigate shared hosting, certificates, and registration patterns, then emitting rules that describe the campaign rather than the page.
Multimodal Approaches and LLM Utilization. LLM and multimodal systems use crawled page content to aid detection or brand attribution, e.g., PhishLLM [7], KnowPhish [14], ScamFerret [15], ChatSpamDetector [16], and ChatPhishDetector [17]. These approaches work best when HTML, text, or screenshots are available. In contrast, PhishLumos tasks LLM agents to query external intelligence, correlate infrastructure clues across URLs, and validate candidate rules, enabling campaign-level detection even when page content is unavailable.
@@ III Proposed System
III-A System Overview
We introduce PhishLumos, an adaptive multi-agent system for detecting advanced phishing attacks, particularly those employing evasive techniques that limit content availability. Figure 2 illustrates the system architecture. Unlike traditional methods that struggle with insufficient data, PhishLumos is designed for such content-inaccessible scenarios. It performs what we term Pivotal Analysis: an adaptive process where LLM-powered agents shift the investigative focus from an evasive URL to its underlying infrastructure footprints. This involves targeted analysis of various external datasets, including historical Domain Name System (DNS) records, Internet Protocol (IP) address information, and Transport Layer Security (TLS) certificate data.
@@ III-A System Overview
Agent Goal / Input / Output
Supervisor Goal: Orchestrate overall analysis flow. Input: URL. Output: Final Report.
URL Analysis Goal: Get historical scan data for a URL.
Input: URL. Output: Scan history, status, IP addresses, brands.
Domain Analysis Goal: Analyze historical DNS/IP data.
Input: Domain Name. Output: IP address history, stability patterns.
IP Analysis Goal: Analyze domains/activity on an IP.
Input: IP Address. Output: Associated domains, past malicious activity.
Certificate Analysis Goal: Analyze TLS certificate history.
Input: Domain Name. Output: Issuer patterns, shared certificates.
Campaign Analysis Goal: Synthesize all findings to profile a campaign.
Input: All collected data. Output: Campaign type, key indicators.
Rule Generation Goal: Generate and validate detection rules.
Input: Key indicators. Output: Validated search queries.
Figure 2: System Architecture of PhishLumos.
@@ III-B Threat Model
Social Dimension. These attacks erode digital equity by targeting vulnerable populations, such as older adults. Beyond financial loss, they shatter the trust required to use essential online services, risking social isolation and loss of autonomy [2].
Technical Dimension. We assume attackers use widely available phishing kits. A key feature of these kits is sophisticated cloaking logic, which delivers malicious content to end-users while presenting benign content to automated scanners. We consider email and Short Message Service (SMS) as primary distribution channels. The analytical focus of PhishLumos is on characterizing and mitigating these coordinated campaigns that deploy many similar, cloaked sites concurrently, rather than isolated, one-off attacks.
@@ III-D Adaptive Pivotal Analysis
At the core of the loop, the LLM assesses the entire Knowledge Base to identify the most promising investigative leads based on a set of triggers. These include: (1) Immediate Triggers, such as significant cloaking behavior (e.g., a newly registered domain redirecting to an unrelated major site like google.com); and (2) Infrastructure Triggers, where accumulated evidence reveals suspicious hosting patterns, such as multiple unrelated domains on a non–Content Delivery Network (CDN) IP address. Importantly, content inaccessibility or cloaking alone does not constitute a positive decision; it only prompts a pivot to infrastructure analysis. The campaign classification requires correlated signals, such as shared IP ranges, TLS certificate reuse, or registrar patterns, before any rule is produced. This mechanism is resilient to domain rotation and short-lived anonymization, because the Supervisor escalates pivots from domains to IP ranges, Autonomous System Numbers (ASN), and certificate issuers when lower-level features become volatile (see Table V). This state-aware, trigger-based orchestration allows PhishLumos to allocate computation where it is most informative, adapting its focus to the tactics used by each campaign.
@@ III-E Data Acquisition and Context Management
PhishLumos employs a systematic process for data acquisition and context management to support its adaptive investigations. External intelligence is sourced programmatically from threat intelligence Application Programming Interfaces (APIs), such as urlscan.io [18] and VirusTotal [19], which provide access to essential datasets such as historical website scans, DNS records, and TLS certificate attributes. Specialized Agents dynamically retrieve this information using dedicated API interaction tools.
@@ III-F Campaign Analysis and Rule Generation
TABLE II: Evidence indicators for campaign classification.
@@
CONFIRMED Consistent brand detection or content hashes, linked by shared IP addresses.
CLOAKED Discrepancy between input and final URL, especially to a known legitimate site.
REUSE Shared IPs/Certificates with historically malicious domains (content may differ).
UNAVAILABLE Inaccessible content (e.g., 404), but infrastructure linked to a known campaign.
UNCLEAR Insufficient correlated evidence to make a confident classification.
@@ IV-A Experimental Setup
Datasets. To evaluate PhishLumos, we constructed two primary datasets: a ground truth collection of phishing URLs and a set of benign URLs for false positive analysis. Our ground truth phishing dataset was built using public data from the Japan Computer Emergency Response Team (JPCERT/CC) [20], spanning from January 2023 to March 2025. JPCERT/CC is Japan’s national Cyber Security Incident Response Team (CSIRT), and its data is highly reliable as human analysts manually verify each reported phishing URL. For each URL, JPCERT/CC publishes the verification date and the targeted brand confirmed by a human analyst; we treat these fields as ground-truth labels. We identified and validated 103 distinct campaigns, comprising 6,020 unique URLs. For false positive analysis, we used a benign dataset of 1,000 safe URLs from the ChatPhishDetector study [17], including legitimate corporate and top-ranking websites [21].
The Challenge of Content-Inaccessible Scenarios. A critical challenge in modern phishing detection is the prevalence of evasive tactics. Our analysis of the ground truth URLs revealed that automated web analysis APIs often fail to retrieve direct phishing content. Cloaking or access errors affected 77.0% of URLs (82.5% of campaigns), preventing direct content retrieval. This rate reflects failures of automated scanners at crawl time; it does not imply that human analysts could not observe phishing content when JPCERT/CC verified the URLs. In PhishLumos, content unavailability is only a trigger to pivot to infrastructure analysis; the system does not assign a positive label on this signal alone. This prevalence of content-inaccessible scenarios underscores the limitations of content-based detectors and highlights the need for infrastructure-focused analysis like PhishLumos. A breakdown reveals diverse evasive tactics: 38.8% used cloaking or deceptive redirection, 37.6% were inaccessible or actively blocked scanners (e.g., with 403/404 errors), and the remaining 23.5% presented ambiguous content with no clear malicious indicators. The problem is escalating, with the rate of these content-inaccessible scenarios increasing from 48.7% in 2023 to 95.4% in 2024, and reaching 100% for the campaigns observed in early 2025.
@@
Implementation and Reproducibility. PhishLumos is implemented in Python 3.11.12, using the LangGraph library (v0.4.5) to manage the multi-agent workflow. All agents use Azure OpenAI’s gpt-4.1-2025-04-14 model. To ensure deterministic outputs and maximize reproducibility, key hyperparameters were set to fixed values. We set a temperature of 0 and a seed of 42 for this purpose. Given the deterministic nature of this setup, each experiment was conducted once per input URL. For external data gathering, the agents interfaced with the public API of urlscan.io [18]. All experiments were run on a MacBook Air with an Apple M2 CPU and 24GB of RAM, running macOS Sequoia 15.5. To facilitate reproducibility, our research assets—source code, full dataset, agent prompts, and execution logs—are available upon request.
Operational model and scalability. In deployment, we invoke the LLM agents only for a small set of seed URLs (e.g., Security Operations Center (SOC) triage). The generated rules compile to lightweight predicates—regular expressions over domains and paths, IP/ASN filters, and TLS issuer fingerprints—that mail gateways and web proxies can evaluate at line rate. This design decouples offline, higher-cost rule synthesis from online, low-cost matching; production systems apply only the rules, not the LLM, to traffic streams.
TABLE III: Overall performance of PhishLumos.
Metric Mean Median
Campaign Coverage (%) 93.0 100.0
New URLs Found (per campaign) 751.4 297.0
Detection Lead Time (h) 529.1 192.8
Execution Time (s) / URL 63.9 60.4
Cost (\$) / URL 0.25 0.23
TABLE IV: Performance comparison in different scenarios.
Scenario Method Precision Recall F1-Score False Positive Rate
@@
TABLE V: Case studies demonstrating PhishLumos’s reasoning. Note: Brand names and technical indicators are anonymized for security. Initial refers to the properties of the input URL, while Final refers to the page attributes after any redirects.
Case Study Generated Rule Insight: Why this rule is intelligent and practical
1. Zero initial scan data Initial Domain: *.co.jp.*.test AND Pivoted from a dead-end URL to its domain, spotting a suspicious domain pattern.
(dead-end URL) Date: [2023-09-02 TO 2024-02-27] This uncovered a regional campaign, finding 405 new URLs.
2. Deceptive cloaking Initial URL: */[unique_string]* AND Identified a non-obvious URL path string as a unique campaign fingerprint,
(redirect to benign site) Date: [2024-11-26 TO 2024-12-24] ignoring the misleading redirect. Discovered 540 new URLs.
3. Advanced cloaking Initial Domain: *.suspicious.test AND Final Domain: official.example Defeated cloaking by flagging the mismatch between the suspicious initial and
(redirect to official site) AND Date: [2023-04-24 TO 2023-05-15] legitimate final domains. Found 281 new cloaked URLs.
4. Scanner blocking Final IP: 203.0.113.43 AND Final Status: 403 Turned a ‘403 Forbidden’ error into a detection rule, treating the evasion tactic
(403 Forbidden error) AND Date: [2024-03-30 TO 2024-07-26] itself as a cloaking signal. Identified 65 new URLs.
5. IP hopping to Final ASN: AS64500 AND Final Domain: *.test Created a resilient rule by targeting the stable ASN instead of transient IPs,
evade blocklists. AND Date: [2023-06-14 TO 2024-01-01] moving up the network stack. Preemptively found 4,118 URLs.
TABLE VI: Ablation study results.